# Construction of 2-D Z-Complementary Array Code Sets with Flexible Lengths for Different System Requirements

### v3

@@ 1. Case I:
    $\begin{split}S_{1}=&\sum_{j=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})},\\ S_{2}=&\sum_{j=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})},\\ S_{3}=&\sum_{j=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})}.\\ \end{split}$ (69)
@@
    1. (1)
      For $S_{1}$, we have $10\leq j\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-$ 1. This implies that $j_{\epsilon}=j^{\prime}_{\epsilon}=0$ as $\tau_{2}<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. Similarly, we can find that $10\leq j_{\xi}\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-$ 1, $\tau_{2}\leq j^{\prime}_{\xi}\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1$. So, using (68), we get
      $\begin{split}S_{1}=&\omega_{q}^{a_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+a_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})}\omega_{q}^{-\left(a_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+a_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right)}\\ &\times\left(\sum_{j_{\xi}=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})}\right)\\ \end{split}$ (71)
@@ 2. Case II:
    $\begin{split}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=\sum_{i=0}^{L_{1}-\tau_{1}-1}\bigg{[}T_{1}+T_{2}\bigg{]},\\ \end{split}$ (76)
@@
    $\begin{split}T_{1}=&\omega_{q}^{a_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+a_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})}\omega_{q}^{-\left(a_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+a_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right)}\\ &\times\left(\sum_{j_{\xi}=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}\right),\\ \end{split}$ (78)
@@
    $\begin{split}T_{2}=&\omega_{q}^{b_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+b_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})}\omega_{q}^{-\left(b_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+b_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right)}\\ &\times\left(\sum_{j_{\xi}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{2}^{1},\mathbf{t}_{2}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}\right).\\ \end{split}$ (79)
@@ 1. Sub-case 1:
      $\begin{split}\sum_{j_{\xi}=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}=&\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}},\\ \sum_{j_{\xi}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{2}^{1},\mathbf{t}_{2}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}=&\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}.\\ \end{split}$ (80)
@@
      $\begin{split}W_{1}(i_{\vartheta})=0\Leftrightarrow&~{}W_{2}(i_{\vartheta})=1,\\ W_{1}(i_{\vartheta})=1\Leftrightarrow&~{}W_{2}(i_{\vartheta})=0.\\ \end{split}$ (81)
@@
      $\begin{split}&\sum_{i=0}^{L_{1}-\tau_{1}-1}\bigg{[}T_{1}+T_{2}\bigg{]}\\ &=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}\sum_{i_{\delta}=0}^{L_{1}-\tau_{1}-1}\bigg{[}\left(\omega_{q}^{a_{1}(\vec{i}_{\delta})-a_{1}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{1}(\vec{i}_{\delta})-b_{1}(\vec{i}^{\prime}_{\delta})}\right)\\ &\quad+\left(\omega_{q}^{a_{2}(\vec{i}_{\delta})-a_{1}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{2}(\vec{i}_{\delta})-b_{1}(\vec{i}^{\prime}_{\delta})}\right)\\ &\quad+\left(\omega_{q}^{a_{2}(\vec{i}_{\delta})-a_{2}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{2}(\vec{i}_{\delta})-b_{2}(\vec{i}^{\prime}_{\delta})}\right)\bigg{]}=0.\end{split}$ (82)
@@ 3. Case III:
    Let $0\leq\tau_{1}<2^{m+1}$ and $-\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau_{2}<0$. This case can be proved similar to Case I.
Combining all these cases, we have the theorem.
Appendix C Proof of Theorem 2
@@
$D^{\nu}_{i,j}=\begin{cases}A^{\nu}_{i,j}(t),&(t-1)(L_{1}+Z_{1}-1)\leq i<\\ &(t-1)(L_{1}+Z_{1}-1)+L_{1},\\ &0\leq j<L_{2},t\in\{1,2,\dots,M\};\\ 0,&\text{otherwise}.\end{cases}$ (86)
This produces $\sum_{i=0}^{N_{1}}\sum_{j=0}^{N_{2}}\left|D^{\nu}_{i,j}\right|^{2}=ML_{1}L_{2}$. Now, we apply cyclic shifts on the 2-D arrays $\mathbf{D}^{\nu}(\nu=1,2,\dots,K)$. We take $Z_{1}$ distinct cyclic shifts (including the 00 shift) in the right direction and $Z_{2}$ distinct cyclic shifts (including the 00 shift) in the downward direction and all their combinations. This procedure produces a new set $\mathcal{E}$ having $KZ_{1}Z_{2}$ distinct 2-D arrays of size $N_{1}\times N_{2}$. We apply Lemma 8 on this set =$\mathcal{E}=\{\mathbf{E}^{\nu}:\nu=1,2,\dots,KZ_{1}Z_{2}-1\}$. Then, we have
$\begin{split}&\overline{K}(\overline{K}\!\!-\!\!1)c^{2}_{max}+\overline{K}\left(ML_{1}L_{2}\right)^{2}\\ &\geq\!\!\!\sum_{s_{i}=0}^{\overline{L}_{1}-1}\sum_{s_{j}=0}^{\overline{L}_{2}-1}\left(\sum_{\nu=0}^{\overline{K}-1}\left|\mathbf{E}^{\nu}_{s_{i},s_{j}}\right|^{2}\right)^{2},\end{split}$ (87)

### v4

@@ 1. Case I:
[…]
@@
    1. 1.
      For $S_{1}$, we have $0\leq j\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1$. This implies that $j_{\epsilon}=j^{\prime}_{\epsilon}=0$ as $\tau_{2}<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. Similarly, we can find that $0\leq j_{\xi}\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1$, $\tau_{2}\leq j^{\prime}_{\xi}\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1$. So, using (59), we get
[…]
@@ 2. Case II:
[…]
@@
[…]
@@
[…]
@@ 1. Sub-case 1:
[…]
@@
[…]
@@
      $\begin{split}\sum_{i=0}^{L_{1}-\tau_{1}-1}\bigg{[}T_{1}+T_{2}\bigg{]}=&\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}\sum_{i_{\delta}=0}^{L_{1}-\tau_{1}-1}\bigg{[}\left(\omega_{q}^{a_{1}(\vec{i}_{\delta})-a_{1}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{1}(\vec{i}_{\delta})-b_{1}(\vec{i}^{\prime}_{\delta})}\right)\\ &+\left(\omega_{q}^{a_{2}(\vec{i}_{\delta})-a_{1}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{2}(\vec{i}_{\delta})-b_{1}(\vec{i}^{\prime}_{\delta})}\right)\\ &+\left(\omega_{q}^{a_{2}(\vec{i}_{\delta})-a_{2}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{2}(\vec{i}_{\delta})-b_{2}(\vec{i}^{\prime}_{\delta})}\right)\bigg{]}\\ &=0,\end{split}$ (73)
@@ 3. Case III:
    Let $0\leq\tau_{1}<2^{m+1}$ and $\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau_{2}<0$. This case can be proved similar to Case I.
Combining all these cases, we have the theorem. ∎
The following example illustrates Theorem 1.
Example 3.
Let $p_{1}=2$, $p_{2}=3$, $q=6$, $\pi_{1}(1)=1$, $\pi_{2}(1)=1$, $a_{\mathbf{s}_{i}^{k},\mathbf{t}_{j}^{k}}^{\bm{\gamma}}$, $f_{1}$, and $f_{2}$ be the same functions as in Example 2. Let the function $f:\mathbb{Z}_{2}^{2}\rightarrow\mathbb{Z}_{6}$ be defined by $f=3x_{1}x_{2}+3$, where $\pi(1)=1,\pi(2)=2$. Then, let the functions $a_{1},b_{1},a_{2},b_{2}:\mathbb{Z}_{2}^{2}\rightarrow\mathbb{Z}_{6}$ be given by $a_{1}=3x_{1}x_{2}+3$, $b_{1}=3x_{1}x_{2}+3+3x_{1}$, and $a_{2},b_{2}$ are generated from Lemma 9. We also take $\bm{\zeta}_{1}=(\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},\mathbf{t}_{1}^{1},\mathbf{t}_{2}^{1})$, where $\mathbf{s}_{1}^{1}=\mathbf{s}_{2}^{1}=(1,1)$ and $\mathbf{t}_{1}^{1}=(0,1)$, $\mathbf{t}_{2}^{1}=(1,2)$ and similarly we take $\bm{\zeta}_{2}=(\mathbf{s}_{1}^{2},\mathbf{s}_{2}^{2},\mathbf{t}_{1}^{2},\mathbf{t}_{2}^{2})$, where $\mathbf{s}_{1}^{2}=\mathbf{s}_{2}^{2}=(1,1)$ and $\mathbf{t}_{1}^{2}=(1,1)$, $\mathbf{t}_{2}^{2}=(1,0)$, so that $\mathbf{t}_{1}^{1},\mathbf{t}_{2}^{1},\mathbf{t}_{1}^{2}$ and $\mathbf{t}_{2}^{2}$ are all distinct. We also take $\vec{d}^{\prime}=(d_{1},d_{2},\tilde{d}_{p})\in\mathbb{Z}_{2}^{2}\times\mathbb{Z}_{3}$, $W_{1},W_{2}$ are generated from (53). Then, we get a 2-D ZCACS of size $12\times 72$ with set size $K=6$. The 2-D AACF sum of 2-D ZCAC $\{\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}}:\bm{\gamma}\in\mathbb{Z}_{2}\times\mathbb{Z}_{3}\}$ is given in the Fig. 1. The size of the rectangular ZCZ is $8\times 6$. The plot of the 2-D ACCF sum of the set $\{\mathcal{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathcal{X}^{\bm{\gamma},\bm{\zeta}_{2}}\}$ is given in Fig. 2.
Figure 1: Plot of 2-D AACF sum corresponding to Example 3
Figure 2: Plot of 2-D ACCF sum corresponding to Example 3
Remark 3.
It should be noted that, fixing $k=1$, and $p_{1}=2$, the set $\mathcal{X}=\{\mathbf{X}^{\gamma}:\gamma\in\mathbb{Z}_{2}\}$ becomes a 2-D ZCAP $\{\mathbf{X}^{0},\mathbf{X}^{1}\}$ of size $2^{m}p\times 2^{m_{1}+1}$ and rectangular ZCZ width is $2^{m+1}\times 2^{m_{1}-1}$. 2-D ZCAP [ChenZCAP, AbhishekZCAP] of this size is reported for the first time in this paper, as a special case.
Corollary 1.
Proof.
$\begin{split}F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}=&\bar{v^{\prime\prime}}\left\{a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}+a_{1}(\vec{d})W_{1}(\tilde{d}_{p})+a_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\}\\ &+v^{\prime\prime}\left\{a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}}+b_{1}(\vec{d})W_{1}(\tilde{d}_{p})+b_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\},\end{split}$ (77)
where $a_{\mathbf{s}_{i},\mathbf{t}_{i}}^{\bm{\gamma}}=R_{\mathbf{t}_{i}}^{\bm{\gamma}}$, i.e., $T_{\mathbf{s}_{i}}=0$ for $i=1,2$. Now, proceeding similarly to the proof of Theorem 1, we get the result. ∎
Corollary 2.
Proof.
4 Analysis of PMEPR Bound
4.1 Row Sequence PMEPR
For a fixed $\bm{\gamma}$ we take the $\mathbb{Q}_{q}$-valued 2-D array $K^{\bm{\gamma}}=\{F_{\tilde{\mathbf{d}}}^{\bm{\gamma},\zeta}:\tilde{\mathbf{d}}\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}\}$. Now, from the proof of Lemma 5, it can be deduced that
$\begin{split}\left|\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}A(K^{\bm{\gamma}}_{m})(\tau_{2})\right|\leq 2\left(\prod_{\alpha=1}^{k}p_{\alpha}-\beta\right)\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}},\end{split}$ (80)
$\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\left|S_{K^{\bm{\gamma}}_{m}}(t)\right|^{2}\leq&\sum_{\beta=1}^{\prod_{\alpha=1}^{k}p_{\alpha}-1}\sum_{\left|\tau_{2}\right|=\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}}\left|\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}A(K^{\bm{\gamma}}_{m})(\tau_{2})\right|\\ &+L_{2}\prod_{\alpha=1}^{k}p_{\alpha}\\ =&2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1}\left(\prod_{\alpha}^{k}p_{\alpha}-1\right)+L_{2}\prod_{\alpha=1}^{k}p_{\alpha}.\end{split}$ (81)
4.2 Column Sequence PMEPR
Similar to the previous case, for the $n$-th column sequence $X_{n}^{col}$, we have
$\begin{split}\left|S_{X_{n}^{col}}(t)\right|^{2}=L_{1}+\sum_{\tau_{1}\neq 0}A(X_{n}^{col})(\tau_{1})\omega_{q}^{-q\tau_{1}\Delta ft}.\end{split}$ (83)
For this, we note that either ${K^{\bm{\gamma}}}^{col}_{n}=\mathcal{I}_{a}$ or ${K^{\bm{\gamma}}}^{col}_{n}=\mathcal{I}_{b}$, where
@@ 4.2 Column Sequence PMEPR
for $(\vec{d},\tilde{d}_{p})\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}$ and $a_{1},a_{2},b_{1},b_{2},W_{1}$ and $W_{2}$ are the functions used in Theorem 1. Note that for $p\neq 2$, we have
$\begin{split}A(\mathcal{I}_{a})(\tau_{1})+A(\mathcal{I}_{b})(\tau_{1})=\begin{cases}(p-\delta)2^{m+1},&\left|\tau_{1}\right|=\delta 2^{m},\delta=0,2,\dots,p-1,\\ 0,&\text{otherwise}.\end{cases}\end{split}$ (85)
For $p=2$, we have
[…]
In that case, $A(\mathcal{I}_{a})(\tau_{1})+A(\mathcal{I}_{b})(\tau_{1})=0,\forall\tau\neq 0$. So, arguing similarly to the previous case, we have
Remark 4.
Take $p=2$, $k=1$ and $p_{1}=2$. Then Theorem 1 provides 2-D ZCAC of array size of the form $2^{n}\times 2^{m}$ and flock size $M=2$, which is comparable to the array size of 2-D GCAPs and 2-D GCASs in [pai2021twodimensional, Th. 12]. The PMEPR of the row and column sequences of 2-D GCAP and 2-D GCAS in [pai2021twodimensional] are upper bounded by $2^{v}$ for some positive integer $v\geq 1$. However, in our construction, for this case, the row and column sequence PMEPRs are upper bounded by $4$ and $2$, respectively. This shows that the proposed 2-D ZCAC emerges as a good candidate to reduce PMEPR compared to [pai2021twodimensional].
Now, we verify the claim made in Remark 4 by an explicit example and simulation results.
Example 4.
We take a 2-D ZCAC from Theorem 1 with the function of (52), where $a_{\mathbf{s_{i}},\mathbf{t_{j}}}^{\gamma}=v_{1}v_{2}+\gamma v_{1}+t_{j}v_{2}+s_{i}v_{3}$, $a_{1}=x_{1}x_{2}+1$, and $a_{2},b_{1},b_{2}$ are derived using Lemma 8 and Lemma 9. Then we get a 2-D ZCAC of size $8\times 16$ and flock size $2$. The row and column sequence PMEPRs are computed to be $3.4271$ and $2$, respectively. In Fig. 3, we have shown the corresponding plot of row and column sequence IAPR and PMEPR bound. We also have taken a binary $8\times 16$-GCAP from [pai2021twodimensional] by using the function $f=(z_{1}z_{2}+z_{2}z_{4}+z_{4}z_{5}+z_{5}z_{3}+z_{3}z_{6}+z_{6}z_{7})+1$ and plotted the corresponding PMEPR/IAPR curve in Fig. 3. It is evident from the figure that column sequence PMEPR of [pai2021twodimensional] is greater than $2$, for this particular example, while the proposed one is always bounded by $2$.
Figure 3: Row and column sequence IAPR/PMEPR
5 Bound on the Set Size
[…]
Lemma 10 ([Roy2024]).
Then, we have
Theorem 2.
[…]
$K\leq\left\lfloor\frac{ML_{1}L_{2}}{Z_{1}Z_{2}}\right\rfloor.$ (91)
Proof.
[…]
$D^{\nu}_{i,j}=\begin{cases}A^{\nu}_{i,j}(t),&(t-1)(L_{1}+Z_{1}-1)\leq i<(t-1)(L_{1}+Z_{1}-1)+L_{1},\\ &0\leq j<L_{2},t\in\{1,2,\dots,M\};\\ 0,&\text{otherwise}.\end{cases}$ (92)
This produces $\sum_{i=0}^{N_{1}}\sum_{j=0}^{N_{2}}\left|D^{\nu}_{i,j}\right|^{2}=ML_{1}L_{2}$. Now, we apply cyclic shifts on the 2-D arrays $\mathbf{D}^{\nu}(\nu=1,2,\dots,K)$. We take $Z_{1}$ distinct cyclic shifts (including the 0 shift) in the right direction and $Z_{2}$ distinct cyclic shifts (including the 0 shift) in the downward direction and all their combinations. This procedure produces a new set $\mathcal{E}$ having $KZ_{1}Z_{2}$ distinct 2-D arrays of size $N_{1}\times N_{2}$. We apply Lemma 10 on this set $\mathcal{E}=\{\mathbf{E}^{\nu}:\nu=1,2,\dots,KZ_{1}Z_{2}-1\}$. Then, we have
$\begin{split}\overline{K}(\overline{K}-1)c^{2}_{max}+\overline{K}\left(ML_{1}L_{2}\right)^{2}&\geq\sum_{s_{i}=0}^{\overline{L}_{1}-1}\sum_{s_{j}=0}^{\overline{L}_{2}-1}\left(\sum_{\nu=0}^{\overline{K}-1}\left|\mathbf{E}^{\nu}_{s_{i},s_{j}}\right|^{2}\right)^{2},\end{split}$ (93)